OpenStudy (anonymous):

Please, PLEASE Help me! Solve the following system of equations. 2x - y + z = -3 2x + 2y + 3z = 2 3x - 3y - z = -4 A. (1, 3, 2) B. (-1, 3, 2) C. (1, -3, 2) D. (1, 3, -2)

OpenStudy (anonymous):

I really need some help with this :l

OpenStudy (anonymous):

Answer is D!

OpenStudy (anonymous):

are you sure??

OpenStudy (anonymous):

Like on a scale of 1-10 how positive are you?

OpenStudy (anonymous):

Well, it's just you go testing Darwin values ​​to see if the equalities will be correct, it looks like I did:   D = (1, 3, -2) 2x - z + y = -3 2 (1) - 3 - 2 = - 3   -3 = -3 -> Correct 2x + 2y = 3z + 2 2 (1) + 2 (3) + 3 (-2) = 2 2 = 2 Correct 3x - 3y - z = -4 3 (1) - 3 (3) +2 = -4 3 +2 = -4 -9 - 4 = - 4 correct

OpenStudy (anonymous):

ops! sorry, it's the google translator! lol